Amoxicillin 25mg/kg/day PO Q12H is ordered for a child with a respiratory infection. Amoxicillin is supplied in an oral suspension of 250 mg/5mL. The child weighs 15.84 lbs.

How many mL should the nurse administer per dose?
Answer #1

Ordered amoxicillin = 25mg/kg/day PO Q12H

PO means orally

Q12H means every 12 hours

The weight of the child= 15.84 lbs= 15.84/2.2= 7.2 kg

As per weight of the child required dose= 7.2X 25= 180 mg/day every 12 hours

per dose requirement= 180/2= 90 mg

Supplied oral suspension = 250 mg/5mL

Means 50 mg/ml

for 90mg required ml= 90/50= 1.8 ml the nurse should administer per dose